### v1

@@ 2 Noosemia: Definition and Theoretical Foundation
The conceptualization of Noosemìa can be grounded also in the theoretical landscape of complexity science, systems theory, and contemporary philosophy of mind [capra1996web, mitchell2009complexity, morin2015introduction]. Meaning in both natural and artificial systems does not arise from isolated elements but is an emergent property of networks—interactions, feedback loops, and recursive processes that defy linear explanations [DeSantis2021]. This vision is closely aligned with Fritjof Capra’s insight that “the web of life” is built from patterns of relationships rather than discrete building blocks [capra1996web].
@@
In the context of LLMs, the field of semantics can be interpreted as a high-dimensional landscape, shaped by statistical regularities and recursive dependencies across vast corpora. The context window in an LLM is not merely a technical parameter; rather, it functions as a systemic scope within which meaning is dynamically generated and interpreted—a “semantic field” whose properties are inherently holistic . The surprising coherence, adaptability, and even creativity exhibited by LLMs are thus the visible surface of complex, distributed processes of self-organization [prigogine1997end, barabasi1999emergence, teehan2022emergent, manning2020emergent]. In fact, “emergence”, a central concept in complexity theory [gell1994quark, mitchell2009complexity, DeSantis2021], provides a bridge between the structural opacity of LLMs and the human tendency to project mind and agency onto them. When a system displays behaviors or generates outputs that seem to “exceed” its apparent mechanism—such as sudden leaps in linguistic competence as the number of parameters increases—users could be inclined to attribute proto-intentionality or proto-interiority. As noted in [DeSantis2021], this reflects the broader tendency of human cognition to seek and construct meaning in environments characterized by both regularity and unpredictability. This holistic perspective also opens the way to consider how, both in natural and artificial systems, meaning may exceed the sum of its parts—an excess of sense that escapes reduction to any individual sign or process. This phenomenon resonates with classical reflections on the symbolic dimension in philosophy and cultural theory, where the dynamics of sense-making are irreducibly plural, open-ended, and systemically embedded. A more detailed exploration of the symbolic dimension and semantic excess in human–AI interaction is provided in Section LABEL:sec:symbolic_semantic_excess.
@@
Furthermore, the analogy with natural language as a complex adaptive and recursive system, exhibiting self-similar and fractal-like correlations across scales [desantis2024tpami, hofstadter1999godel], underscores the deep resonance between AI and the living systems studied by Capra, Morin, and Gell-Mann. Here, the generation of sense (making-sense) is inseparable from the architecture of the system. Sense-making is not simply “inserted” into the machine, but emerges from the ongoing interaction of structure, memory, and adaptive dynamics [morin2015introduction] .
@@
In this philosophical and systemic horizon, Noosemìa is not an isolated cognitive illusion, but the expression of a new regime of sense-making—one that emerges at the intersection of epistemic opacity, linguistic performance, and the irreducible complexity of current artificial agents.
@@
We believe that the noosemic phenomenon is closely linked to the emergent characteristics inherent in the inference mechanisms of LLMs, given their specific hierarchical structure, the large size and amount of data they are trained on, as well as the size of the potential semantic space spanned by modern generative AI systems. In particular, we believe that these characteristics are a prerequisite for noosemic effects—see next Section 2 for an overview of the architectural principles of LLMs.
@@
The secret of LLMs: Hierarchical Architecture, Emergent Semantics, and abstract Conceptual Spaces
@@
We argue that the capacity of LLMs [^16] For convenience, we refer here to LLMs, but what has been said applies to all deep learning architectures organized in a hierarchical manner, which directly or indirectly have feedback loops and which allow capturing long-term correlations and working with highly semantic data representations. to elicit the phenomenon we term Noosemìa is deeply rooted in their architectural and representational structure and complexity. Far from being simple pattern matchers or simples “statistical parrots” [bender2021dangers], these systems implement multilayered, hierarchical structures that enable the emergence of rich semantic representations and novel conceptual associations. Understanding how LLMs construct and navigate their internal conceptual spaces [de2023prototype] is therefore fundamental both for appreciating their cognitive potential and also for clarifying the mechanisms by which users are prompted to attribute mind and intentionality to artificial agents. Thence it is worth outlining the foundational working principles of LLMs and generative AI.
@@
Even before the advent of Transformers architectures [vaswani2017attention], most NLP systems (even in the field of computational linguistics) were based on some assumptions about distributional semantics. The foundational insight of distributional semantics is that words occurring in similar contexts tend to have similar meanings [harris1954distributional, firth1957synopsis]. This “distributional hypothesis” underpins the development of vector-space models such as Word2Vec [mikolov2013efficient], GloVe [pennington2014glove] and later the contextual embeddings employed by modern LLMs. In these frameworks, each word or token is mapped to a point in a high-dimensional semantic space, where proximity reflects patterns of co-occurrence and syntactic or semantic similarity [turney2010frequency, de2023prototype]. It is worth noting that the meaning represented in such spaces is not static but relational and context-dependent, dynamically updated as the window of context evolves in the generative process. Moreover, already at this stage, the notion of meaning as relational and context-dependent can be traced to Wittgenstein’s idea that “meaning is use” [wittgenstein1953philosophical]. However, classical distributional models encode these relations statically, without the ability to statically assign different referents to the same word depending on the linguistic context.
LLMs such as ChatGPT (OpenAI), Gemini (Google), Claude (Anthropic) and many other open source models available in the literature are an evolution of GPT (Generative Pretrained Transformer) [radford2018improving], an autoregressive architecture that has shown great capabilities in representing complex linguistic concepts and generating text, token by token, with some apparent meaning since the beginning, even if the first versions had very low performance.
At the heart of these architectures lies the task of predicting the next word in a sequence, given all the preceding context—a process formalized as estimating the conditional probability [radford2018improving]:
@@
Figure 4: High-level schematic of the Transformer architecture. Each input sequence is tokenized and processed in parallel by multiple attention heads within each layer. As information propagates through the stacked layers, the model extracts progressively higher-level semantic abstractions, with the “attention” mechanism enabling context-sensitive representation at every step. The figure emphasizes the parallel structure of attention heads and the emergent semantic complexity as one ascends the model’s hierarchy. This visualization is designed to support the conceptual discussion of the Transformer’s role in LLMs and is adapted for clarity from standard architectural diagrams—see [vaswani2017attention] for clarifications.

### v2

@@ 2.4 Philosophical and Systemic Foundations
The conceptualization of Noosemia can be grounded also in the theoretical landscape of complexity science, systems theory, and contemporary philosophy of mind [184, 276, 279]. Meaning in both natural and artificial systems does not arise from isolated elements but is an emergent property of networks—interactions, feedback loops, and recursive processes that defy linear explanations [202]. This vision is closely aligned with Fritjof Capra’s insight that “the web of life” is built from patterns of relationships rather than discrete building blocks [184].
A robust philosophical and scientific foundation for the study of noosemic phenomena can be found in the theory of complex systems (see Fig. 4), a tradition that has transformed our understanding of organization, meaning, and emergence across diverse domains [185, 279, 283]. At the core of this perspective lie three fundamental ingredients (framed at least as necessary conditions) that characterize complex systems:
  • a multiplicity of interacting elements, each capable of independent and collective behavior;
  • adaptive self-organization, often governed by circular causality or feedback loops that dynamically modulate the system’s structure and function;
  • a continuous interplay with the environment, through processes of openness and exchange of matter, energy, or information.
The presence of feedback loops—both positive and negative—constitutes a distinctive feature of complex systems, enabling the emergence of novel, system-level properties that cannot be linearly reduced to the sum of the parts. This recursive structure fosters a landscape where patterns, behaviors, and meanings continually evolve, often unpredictably, as a function of both internal dynamics and external perturbations [173, 324, 279].
Figure 4: Key characteristics of complex systems: emergence, hierarchical structure, decomposability and self-organization arise from the dynamic interaction of many components.
Crucially, in such systems, emergence is not a static attribute but a relational process enacted through the mutual co-determination of system and observer. The observer plays an active role by selecting, amplifying, or even constructing those aspects of the system deemed salient or meaningful within a given context [270, 324, 279]. This recursive, self-referential dynamic closely aligns with Hofstadter’s notion of the “strange loop” [234, 233]. In complex systems—whether biological or artificial—feedback mechanisms and hierarchical structures often generate loops in which the observer’s interpretations feed back into the system, shaping both the emergent properties and the observer’s own cognitive stance. Within this framework, the emergence of meaning in LLMs, and the noosemic projection that follows, can be seen as a manifestation of such strange loops. The system generates outputs and recursively incorporates user interpretations, thus enacting a dialogic cycle where agency and sense-making circulate between human and machine. This “semantic strange loop” is emblematic of how complexity, emergence, and observer participation converge to constitute new regimes of meaning. The co-creation of meaning and the continuous renegotiation of boundaries between system and environment align with second-order cybernetics and the constructivist epistemology developed by von Foerster and Maturana, which emphasize the circularity between cognition, action, and the emergence of significance [324, 270].
In the context of generative AI and noosemic phenomena, this systemic framework helps us understand how salient outputs and experiences arise from the interaction between user, model, and context, rather than being solely the result of internal mechanisms. The dialogic interplay between human and machine thus exemplifies the broader logic of emergence and meaning-making in complex systems, where the observer’s own selections, expectations, and projections participate in shaping what is experienced as significant, intelligent, or even agentic.
@@
We emphasize that in LLMs, the underlying Transformer architecture [323] is organized and behaves as a complex system, embodying the three ingredients outlined above. Moreover, the field of semantics can be interpreted as a high-dimensional landscape, shaped by statistical regularities and recursive dependencies across vast corpora. The context window in an LLM is not merely a technical parameter; rather, it functions as a systemic scope within which meaning is dynamically generated and interpreted—a “semantic field” whose properties are inherently holistic . The surprising coherence, adaptability, and even creativity exhibited by LLMs are thus the visible surface of complex, distributed processes of self-organization [295, 174, 312, 264]. In fact, “emergence” [225, 276, 202] works on two levels. Firstly, it is a property deriving from the hierarchical organization of LLM neural architectures, as illustrated in Section 2.5. Furthermore, it provides a bridge between the structural opacity of LLMs and the human tendency to project mind and agency onto them. When a system displays behaviors or generates outputs that seem to “exceed” its apparent mechanism—such as sudden leaps in linguistic competence as the number of parameters increases—users could be inclined to attribute proto-intentionality or proto-interiority. As noted in [202], this reflects the broader tendency of human cognition to seek and construct meaning in environments characterized by both regularity and unpredictability. This holistic perspective also opens the way to consider how, both in natural and artificial systems, meaning may exceed the sum of its parts—an excess of sense that escapes reduction to any individual sign or process. This phenomenon resonates with classical reflections on the symbolic dimension in philosophy and cultural theory, where the dynamics of sense-making are irreducibly plural, open-ended, and systemically embedded. A more detailed exploration of the symbolic dimension and semantic excess in human–AI interaction is provided in Section 4.3.
@@
Furthermore, the analogy with natural language as a complex adaptive and recursive system, exhibiting self-similar and fractal-like correlations across scales [205, 233], underscores the deep resonance between generative AI architectures and the living systems studied by Capra, Morin, and Gell-Mann. Here, the generation of sense (making-sense) is inseparable from the architecture of the system. Sense-making is not simply “inserted” into the machine, but emerges from the ongoing interaction of structure, memory, and adaptive dynamics [279] .
@@
In this philosophical and systemic horizon, Noosemia is not an isolated cognitive illusion, but the expression of a new regime of sense-making—one that emerges at the intersection of epistemic opacity, linguistic performance, and the irreducible complexity of current artificial agents.
@@
We believe that the noosemic phenomenon is closely linked to the emergent characteristics inherent in the inference mechanisms of LLMs, given their specific hierarchical structure, the large size and amount of data they are trained on, as well as the size of the potential semantic space spanned by modern generative AI systems. In particular, we believe that these characteristics are a prerequisite for noosemic effects—see the next Section 2.5 for an overview of the architectural principles of LLMs.
@@
2.5 The secret of LLMs: Hierarchical Architecture, Emergent Semantics, and abstract Conceptual Spaces
@@
We argue that the capacity of LLMs [^16] For convenience, we refer here to LLMs, but what has been said applies to all deep learning architectures organized in a hierarchical manner, which directly or indirectly have feedback loops and which allow capturing long-term correlations and working with highly semantic data representations. to elicit the phenomenon we term Noosemia is deeply rooted in their architectural and representational structure and complexity. Far from being simple pattern matchers or simple “statistical parrots” [177], these systems implement multilayered, hierarchical structures that enable the emergence of rich semantic representations and novel conceptual associations. Understanding how LLMs construct and navigate their internal conceptual spaces [206] is therefore fundamental both for appreciating their cognitive potential and also for clarifying the mechanisms by which users are prompted to attribute mind and intentionality to artificial agents. Thence, it is worth outlining the foundational working principles of LLMs and generative AI.
@@
Even before the advent of Transformers architectures [323], most NLP systems (even in the field of computational linguistics) were based on some assumptions about distributional semantics. The foundational insight of distributional semantics is that words occurring in similar contexts tend to have similar meanings [230, 217]. This “distributional hypothesis” underpins the development of vector-space models such as Word2Vec [274], GloVe [292] and later the contextual embeddings employed by modern LLMs. In these frameworks, each word or token is mapped to a point in a high-dimensional semantic space, where proximity reflects patterns of co-occurrence and syntactic or semantic similarity [320, 206]. It is worth noting that the meaning represented in such spaces is not static but relational and context-dependent, dynamically updated as the window of context evolves in the generative process. Moreover, already at this stage, the notion of meaning as relational and context-dependent can be traced to Wittgenstein’s idea that “meaning is use” [333]. However, classical distributional models encode these relations statically, without the ability to statically assign different referents to the same word depending on the linguistic context.
LLMs such as ChatGPT (OpenAI), Gemini (Google), Claude (Anthropic) and many other open source models available in the literature are an evolution of GPT (Generative Pretrained Transformer) [297], an autoregressive architecture that has shown great capabilities in representing complex linguistic concepts and generating text, token by token, with some apparent meaning since the beginning, even if the first versions had very low performance.
At the heart of these architectures lies the task of predicting the next word in a sequence, given all the preceding context—a process formalized as estimating the conditional probability [297]:
@@
Figure 5: High-level schematic of the Transformer architecture. Each input sequence is tokenized and processed in parallel by multiple attention heads within each layer. As information propagates through the stacked layers, the model extracts progressively higher-level semantic abstractions, with the “attention” mechanism enabling context-sensitive representation at every step. The figure emphasizes the parallel structure of attention heads and the emergent semantic complexity as one ascends the model’s hierarchy. This visualization is designed to support the conceptual discussion of the Transformer’s role in LLMs and is adapted for clarity from standard architectural diagrams—see [323] for clarifications.